An unsuspecting bird is coasting along in an easterly direction at 1.00 mph when a strong wind from the south imparts a constant acceleration of 0.300 m/s2. If the wind\'s acceleration lasts for 3.40 s, find the magnitude r and direction θ (measured counterclockwise from the easterly direction) of the bird\'s displacement over this time interval. (HINT: assume the bird is originally travelling in the x direction and there are 1609 m in 1 mile.)An unsuspecting bird is coasting along in an easterly direction at 1.00 mph when a strong wind from theMapD south imparts a c
Answer #1

Assume that

+\small \hat{i} = +x = East

+\small \hat{j} = +y = North

Part 1)

vi = 1.00 mph East = 1.00 \small \hat{i} + 0\small \hat{j} mph

Convert mph to m/s

  vi = (1.00\small \hat{i} + 0 \small \hat{j} ) mph x (1609 / 1 mile) x (1 hr / 3600 s) = 0.447 \small \hat{i} + 0 \small \hat{j} m/s

a = 0.300 m/s2 From the South = 0.300 m/s2 North = 0 \small \hat{i} + 0.300 \small \hat{j} m/s2

Equation of motion:

2 xf ro+vot at

u = 0

  (0.447 +0x3.40s +x (0+0.300jm/s(3.40s)2

  (1.519i0j)(02+1.734)

r1.5191.734

rf V1.51921.7342 2.30m

r = 2.30 m

Direction (1.734 1.519 48.789 0 tan North of East or 41.22o compass

Part 2

  vi = (0.447 \small \hat{i} + 0 \small \hat{j} ) m/s

a = 0.500 m/s2 34.0 deg above or below East (x-axis, y-axis) = 0.500 m/s2 x (cos(34.0) \small \hat{i} + sin (34.0) \small \hat{j} ) =

a = (0.414\small \hat{i} + 0.279 \small \hat{j} ) m/s2

2 xf vot at

  r (0.447 0) x 3.40s(0.414i+0.279) x 3.40s

\small x_f=(1.519\hat{i}+0\hat{j})+(0.703\hat{i}+0.474\hat{j})

(1.519i0.703)(00.474)2.222i+0.474

F2.2221 0.474j

\small \theta =\tan^{-1}\left ( \frac{0.474}{2.222} \right )=12.04^{\circ} cartesian or 12.04o North of East or 77.96o Compass